The molecular weight of oxygen and nitrogen are 32 and 28 respectively At 15 degree celcius the presence of 1 gm oxygen be same as that of 1 gm nitrogen in the same bottle at the temperature

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

pressure p=nRT/V

volume is same for both the gases

number of mole for oxygen = 1/32

number of mole for nitrogen = 1/28

given presuure is same

 n1RT1/V1  = n2RT2/V2

 (1/32) R (273+15)/V  =  (1/28) R T /V

  T=28/32 (288)=252 K

   =-21 C

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin